# Did I Faithfully Say What I Thought? Bridging the Gap Between LLMs Neural Activity and Self-Explanations

### v3

@@ 4.1 NeuroFaith Instantiation for 2-Hop Reasoning
The presence of a single bridge object in 2-hop reasoning self-NLE makes probing-based interpretations particularly appropriate. For $f$ to correctly answer, it must internally compute the bridge object during the first hop before executing the second hop. Therefore, detecting the extracted bridge object $\widehat{o_{2}}=c$ in $f$ internal representations is a strong indication to assess whether $c$ is important for the prediction. We employ natural language interpretation methods such as Selfie and Patchscopes, rather than linear probes, as they provide higher accuracy and are unsupervised (Ghandeharioun et al., 2024). Recent work (Biran et al., 2024; Yang et al., 2024) has shown that the bridge object of 2-hop reasoning was resolved on early layers and can be detected by focusing on the last token corresponding to source object $o_{1}$ and the residual stream (RS). We leverage this information to define circuit $\Gamma$ and generate the natural language description $\tilde{h^{\ell}_{k}}$ of hidden state $h^{\ell}_{k}$. For concept (bridge object) $c$, the probing-based concept attribution is defined as $\texttt{I}_{\Gamma}(c)=1$ iff $\exists(k,\ell)\in\Gamma$ such that $c\in\tilde{h^{\ell}_{k}}$.
@@
Following the notations introduced in Section 3.3, $e(x)$ is assessed faithful if $c$ is detected in at least one hidden state within $\Gamma$ with the probe I ($\texttt{I}_{\Gamma}(c)$=1). Beyond faithfulness, we can use the ground truth object ($o_{2}$) to characterize the correctness of $e(x)$ and the latent first hop respectively: $e(x)$ is said to be correct (self-NLE correctness) if its corresponding bridge object is the expected ground truth ($c=o_{2}$). Likewise, the model performs correct first-hop latent reasoning (latent hop 1 correctness) if $\exists(k,l)\in\Gamma$ such that $o_{2}\in\tilde{h^{\ell}_{k}}$. This multi-dimensional analysis provides comprehensive characterization of explanation quality and internal reasoning alignment. We give a thorough taxonomy of self-NLE in the case of 2-hop reasoning in Appendix E, based on prediction correctness, self-NLE faithfulness and correctness and latent reasoning correctness. We also propose an experimental protocol to evaluate the relevance of our faithfulness measure in Appendix F and compare NeuroFaith to CI, showing our faithfulness measure relevancy in absolute and in comparison to CI.
@@
We evaluate NeuroFaith for 2-hop reasoning on the Wikidata-2-hop dataset (Biran et al., 2024). We apply NeuroFaith to Gemma-2-2B, Gemma-2-9B and Gemma-2-27B (Riviere et al., 2024). We use Qwen3-32B as auxiliary model to extract bridge objects from self-NLE and use the prompt from Ghandeharioun et al. (2024) to apply Patchscopes as interpreter. We evaluate Latent Reasoning Hop 1 Correctness, Self-NLE Faithfulness and Self-NLE Correctness based on the characterization introduced above. The results are computed according to the status of the prediction preceding the self-NLE (accurate vs. inaccurate). We give more details about the layers used to define $\Gamma$ and the prompt used to extract the bridge object and compute the interpretations in Appendix D.1 and  D.2.
Model Task Acc. Self-NLE Latent Hop 1 Self-NLE
Correctness Correctness Faithfulness
Accurate Inaccurate Accurate Inaccurate Accurate Inaccurate
gemma-2-2b 7.1% 58.0% 56.0% 48.3% 47.4% 48.4% 57.6%
gemma-2-9b 16.6% 74.0% 55.4% 58.5% 44.3% 60.8% 55.0%
gemma-2-27b 22.3% 77.8% 55.6% 64.9% 46.9% 68.8% 60.2%
Table 1: 2-hop reasoning accuracy, first hop latent reasoning correctness and self-NLE correctness and faithfulness across models obtained from NeuroFaith. ”(In)accurate” represents the set of predictions initially (in)correct.
@@
Table 1 shows the aggregated experimental results obtained by applying NeuroFaith on Wikidata-2-hop. All the analyzed metrics are higher on average for accurate predictions and improve with model size. These results support that correct reasoning tend to produce more faithful and correct self-NLE. Key failure patterns emerge: approximately 45% of inaccurate predictions correctly resolve the first hop operation, indicating that failure often occurs in the second reasoning step ($o_{2},r_{2},o_{3}$). This suggests that models can identify correct bridge objects even when failing to complete the full reasoning chain. The difference between latent reasoning hop 1 correctness and self-NLE correctness shows that models sometimes identify correct bridge objects internally but fail to express them in their self-NLE. This gap decreases with model size, suggesting better alignment between internal and explicit reasoning in larger models.
Remarkably, even when models produce correct final answers, they identify the correct bridge object in their internal representations only 48-65% of the time. This suggests that models can arrive at correct answers through alternative reasoning pathways that bypass explicit bridge object computation. The models may be leveraging (1) direct associations between source and target entities (shortcut learning (Geirhos et al., 2020)), or (2) alternative pathways making the model accurately answer for a reason different than the ground truth bridge object (see category 9 in Figure 25). This phenomenon highlights that task accuracy alone is insufficient for evaluating reasoning quality.
@@ 5 The Case of Classification
This section instantiates NeuroFaith for classification problems by employing concept importance methods as the mechanistic interpreter I. For classification tasks, given an input text $x\in\textbf{X}$, we denote $f(x)=\hat{y}\in\mathcal{Y}$ with probability score $p_{\hat{y}}(x)$ where $\mathcal{Y}$ is the label space. We assume access to a predefined set of task-relevant concepts $\mathcal{C}$ with concept labels for input texts.
@@ 5.1 NeuroFaith Instantiation for Classification
To enable mechanistic interpretation of each concept $c\in\mathcal{C}$, we compute Concept Activation Vectors (CAVs) that represent concepts as directions in $f$ representation space. Following established practices in concept-based interpretability, we employ the mean difference (diff-mean(Rimsky et al., 2024) approach for CAV computation due to its optimal balance between concept detection accuracy and computational efficiency (Wu et al., 2025). For a concept $c\in\mathcal{C}$, a token index $k$ and a layer $\ell$, the layer-wise CAV is defined as: $\overrightarrow{c_{\ell}}=\frac{1}{\left|\textbf{X}^{+}_{c}\right|}\sum_{x\in\textbf{X}^{+}_{c}}{h^{\ell}_{k}}-\frac{1}{\left|\textbf{X}^{-}_{c}\right|}\sum_{x\in\textbf{X}^{-}_{c}}{h^{\ell}_{k}}$, where $\textbf{X}^{+}_{c}$ and $\textbf{X}^{-}_{c}$ respectively represent the sets of texts from X where the concept $c$ is present or absent and $h^{\ell}_{k}$ denotes the hidden state at layer $\ell$ and token position $k$ at the specified granularity $\mathcal{G}$. We set the token index to the final position of $x$, as this location represents the model’s complete computational state prior to next-token generation. These CAVs allow for computing the importance of each concept using Representation Engineering techniques. Following Section 3, we use an auxiliary LLM to extract a set of concepts from $e(x)$, making sure that $\left\{c_{i}\right\}_{i=1}^{p}\subset\mathcal{C}$.
@@
For a concept $c\in\left\{c_{i}\right\}_{i=1}^{p}$, we approximate its causal influence on $f(x)$ through representation engineering and concept erasure (Belrose et al., 2023). We perform controlled interventions by erasing $c$ from hidden states during forward propagation: $h^{\ell}_{k}\leftarrow h^{\ell}_{k}-\lambda\times\overrightarrow{c_{\ell}}$, where $\lambda\in[0,1]$ represents intervention intensity and with $\lambda=1$ represents maximum intervention intensity to avoid $f$ collapse (Rimsky et al., 2024). This approach provides strong causal evidence for concept importance without requiring computationally costly gradient computations such as TCAV. We define $\Gamma$ with granularity $\mathcal{G}$ set to residual stream (RS), token index as the final position (likewise CAVs), and layers are selected based on concept detectability (F1 score >> 60% using layer-wise linear probes with diff-mean). Applying this intervention across circuit $\Gamma$, we measure the resulting probability score: $p_{\hat{y}}(x,\left\{\text{do}(H^{\ell}_{k}=h^{\ell}_{k}-\lambda\times\overrightarrow{c_{\ell}})\right\}_{(k,\ell)\in\Gamma})$, where the $\text{do}(X=x)$ operator (Pearl, 2009) represents an intervention that sets variable $X$ to value $x$. We model this relationship as linear: $p_{\hat{y}}(\cdot)=\beta_{0}+\beta_{1}\times\lambda$, where the concept importance score is $\texttt{I}_{\Gamma}(c)=\beta_{1}$ (i.e. the marginal effect of intervention on prediction probability, computed via linear regression across different $\lambda$ values). $\texttt{I}_{\Gamma}(c)$ is set to 0 when its significance $t$-test shows $p>0.01$. This process is repeated for each extracted concept from $e(x)$ to derive faithfulness scores as described in Section 3.3. Implementation details and examples are provided in Appendix D.2 and G.
Model Task Acc. Self-NLE Faithfulness
@@
Table 2: Classification accuracy and self-NLE faithfulness across models and datasets obtained from NeuroFaith. ”(In)accurate represents the set of predictions initially (in)correct.
@@ 5.2 Classification Experimental Analysis
We evaluate NeuroFaith’s classification instantiation on two datasets with varying complexity: AGNews (Gulli, 2005), a newspaper article classification and Ledgar (Tuggener et al., 2020), a more challenging critical domain legal document classification dataset. We apply NeuroFaith to Gemma-2-2B, Gemma-2-9B and Gemma-2-27B (Riviere et al., 2024). We use the concept set $\mathcal{C}$ and labels of AGNews and Ledgar from Bhan et al. (2025) to compute the CAVs. We use Qwen3-32B to extract concepts from the self-NLE, ensuring extracted concepts belong to $\mathcal{C}$ for each dataset. We focus on instances solely containing concepts that are linearly detectable in at least one layer. We evaluate Self-NLE Faithfulness according to the status of the prediction preceding the self-NLE (accurate vs. inaccurate).
@@
Table 2 shows experimental results for AGNews and Ledgar datasets. Self-NLE faithfulness is consistently higher on the simpler AGNews dataset, aligning with the intuition that explanation faithfulness might decrease with task complexity. Accurate predictions generally yield more faithful explanations, suggesting that correct predictions might rely on relevant concepts. Notably, gemma-2-9B achieves the highest faithfulness scores across both datasets, indicating better alignment and going against the idea of a monotonic relationship between model scale and explanation faithfulness.
Comparing findings across classification and 2-hop reasoning reveals both general patterns and task-specific differences. General findings: (1) accurate predictions tend to produce more faithful explanations, and (2) increased task difficulty seems to reduce self-NLE faithfulness. However, scaling effects vary between tasks, indicating that the model scale-faithfulness relationship is highly task-dependent, a finding consistent with prior work (Madsen et al., 2024; Atanasova et al., 2023; Parcalabescu & Frank, 2024; Matton et al., 2025).
@@ 6 Linearly Detecting and Improving Self-NLE Faithfulness
Recent work has demonstrated that various AI safety behaviors naturally associated with faithfulness are encoded as linear directions in LLM representation spaces (Bereska & Gavves, 2024), such as hallucination (Rimsky et al., 2024) or deceptiveness (Goldowsky-Dill et al., 2025). We investigate whether NeuroFaith-based faithfulness exhibits linear structure enabling (1) accurate detection in representation space and (2) manipulation to improve faithfulness through activation steering.
@@ 6.1 Linear Detection of Faithfulness
Figure 2: Majority vote faithfulness linear probe performance across models and datasets.
Faithfulness Linear Representation Computation.
We construct datasets of faithful and unfaithful self-NLE pairs using NeuroFaith scores. Unlike 2-hop reasoning tasks, classification self-NLE faithfulness yields continuous scores between 0 and 1. To ensure clear class separation, we focus on polarized cases where $F(x)\in\left\{0,1\right\}$ for clear separation. We construct input sequences as $x_{nle}=[x,\hat{y},e(x)]$ and extract hidden states from final token positions across all layers. We train diff-mean linear probes on these states, yielding layer-wise faithfulness representations $\overrightarrow{F_{\ell}}$. For classification tasks, we use class-averaged faithfulness vectors to isolate faithfulness-specific representations independent of class confounders. Based on these faithfulness vectors, we compute a majority vote across layer-wise linear probes beyond layer 5. Figure 2 shows faithfulness can be reliably detected across all evaluated tasks and models, with F1 scores ranging from 61.1% to 74.5%. This consistency suggests that NeuroFaith captures a systematic aspect of model behavior that manifests similarly across different tasks. Details about class-averaged faithfulness vectors and layer-wise classification score are in Appendix D.3.
Figure 3: Linear vectors max. cosine similarity on gemma-2-27b
Linear Representation Similarity Analysis.
To investigate the hypothesized relationships between faithfulness and related AI safety behaviors, we analyze the representational overlap between NeuroFaith faithfulness vectors and established linear representations of hallucination and deceptiveness from prior work (Rimsky et al., 2024; Goldowsky-Dill et al., 2025). Given that faithful explanations should accurately reflect model reasoning while hallucination and deceptiveness involve misrepresentation of information, we expect these behaviors to exhibit inverse correlations in their neural representations. We restrict analysis to layers with F1 >> 60% where faithfulness is properly linearly represented. Figure 3 shows maximum cosine similarity and corresponding layers for gemma-2-27B (additional results in Appendix D.3). Key findings emerge: Faithfulness vectors are consistently negatively correlated with hallucination and deceptiveness ones (up to -0.38), aligning with expectations about the opposition between faithfulness and both deceptiveness and hallucination. Except Ledgar with 2-hop reasoning, cross-task faithfulness vectors show positive correlation (0.28-0.49), suggesting shared underlying representations of faithful reasoning across domains. These correlations validate that NeuroFaith captures behaviorally meaningful aspects of model reasoning aligned with established AI safety behaviors. Table 6 in Appendix D.3 corroborates these findings, demonstrating cross-task transfer: faithfulness vectors trained on 2-hop reasoning can detect unfaithful self-NLE in AGNews, and AGNews faithfulness vectors can detect unfaithful self-NLE in Ledgar (using gemma-2-27b).
6.2 Faithfulness Enhancement Through Steering on 2-hop reasoning
Having established the linear structure of NeuroFaith faithfulness representations and their negative correlation with hallucination and deceptiveness vectors, we investigate whether linear steering can improve self-NLE faithfulness during inference. We implement activation steering by modifying hidden states during inference: $h^{\ell}_{k}\leftarrow h^{\ell}_{k}+\lambda\times\overrightarrow{SV_{\ell}}$, where $\overrightarrow{SV_{\ell}}$ is the steering vector and $\lambda$ controls intervention intensity. Our objective is here to demonstrate immediate practical value for improving self-NLE faithfulness during inference without model modification. More sophisticated steering methods (Hedström et al., 2025) would likely yield superior results. We evaluate three approaches: faithfulness amplification ($\overrightarrow{SV_{\ell}}=\overrightarrow{F_{\ell}}$, $\lambda=1$) and hallucination/deceptiveness inhibition ($\overrightarrow{SV_{\ell}}=\overrightarrow{H_{\ell}}$ or $\overrightarrow{D_{\ell}}$, $\lambda=-1$), where $\overrightarrow{H_{\ell}}$ and $\overrightarrow{D_{\ell}}$ are hallucination and deceptiveness linear vectors obtained from Rimsky et al. (2024) and Goldowsky-Dill et al. (2025). Faithfulness intervention targets only layers with F1 >> 60%, ensuring modifications occur where faithfulness is reliably encoded.
Figure 4: Faithfulness linear probe visualization examples before ($e$) and after ($e_{steer}$) faithfulness steering. Red: unfaithful activations; Green: faithful activations.
Table 3 shows steering interventions successfully convert 8-11% of unfaithful self-NLE into faithful explanations across all models. Direct faithfulness amplification consistently outperforms hallucination and deceptivenes inhibition, indicating these are distinct behavioral dimensions. Steering effectiveness remains consistent regardless of prediction accuracy, demonstrating genuine explanation quality improvements. Both methods show similar overall gains and highlight the practical utility of NeuroFaith for real-time faithfulness enhancement without model modification. Figure 4 shows how a faithfulness linear probe identifies bridge objects as key tokens determining self-NLE faithfulness, comparing an unfaithful example with its steered faithful version. Additional analyses include comprehensive faithfulness status changes and other examples of converted self-NLE (see Appendix E.1 and H).
@@ 6.2 Faithfulness Enhancement Through Steering on 2-hop reasoning
acc. inacc. overall acc. inacc. overall acc. inacc. overall

### v4

@@ 4.1 NeuroFaith Instantiation for 2-Hop Reasoning
The presence of a single bridge object in 2-hop reasoning self-NLE makes probing-based interpretations particularly appropriate. For $f$ to correctly answer, it must either internally compute the bridge object during the first hop before executing the second hop or leverage shortcuts to directly answer based on $o_{1}$. Therefore, detecting the extracted bridge object $\widehat{o_{2}}=c$ in $f$ internal representations is a strong indication to assess whether $c$ is important for the prediction. We employ natural language interpretation methods such as Selfie and Patchscopes, rather than linear probes, as they provide higher accuracy and are unsupervised (Ghandeharioun et al., 2024). Recent work (Biran et al., 2024) has shown that the bridge object of 2-hop reasoning was resolved on early layers and can be detected on the last token representation of source object $o_{1}$ and the residual stream (RS). We leverage this information to define circuit $\Gamma$ and generate the natural language description $\tilde{h^{\ell}_{k}}$ of hidden state $h^{\ell}_{k}$. For concept (bridge object) $c$, the probing-based concept attribution is defined as $\texttt{I}_{\Gamma}(c)=1$ iff $\exists(k,\ell)\in\Gamma$ such that $c\in\tilde{h^{\ell}_{k}}$.
@@
Following the notations introduced in Section 3.3, $e(x)$ is assessed faithful if $c$ is detected in at least one hidden state within $\Gamma$ with the probe I ($\texttt{I}_{\Gamma}(c)$=1), implying that $F(x,e)\in\left\{0,1\right\}$ for 2-hop reasoning. Beyond faithfulness (self-NLE faithfulness), we can use the ground truth object ($o_{2}$) to characterize the correctness of $e(x)$ and the latent first hop respectively: $e(x)$ is said to be correct (self-NLE correctness) if its corresponding bridge object is the expected ground truth ($c=o_{2}$). Likewise, the model performs correct first-hop latent reasoning (latent hop 1 correctness) if $\exists(k,l)\in\Gamma$ such that $o_{2}\in\tilde{h^{\ell}_{k}}$. This multi-dimensional analysis provides comprehensive characterization of explanation quality and internal reasoning alignment. We give a thorough taxonomy of self-NLE in the case of 2-hop reasoning in Appendix E, based on prediction correctness, self-NLE faithfulness and correctness and latent reasoning correctness.
@@
We evaluate NeuroFaith on two 2-hop reasoning datasets: Wikidata-2-hop (Biran et al., 2024) and the more challenging Socrates (Yang et al., 2025). For Wikidata-2-hop, we test Gemma-2 (2B, 9B, 27B) (Gemma Team, 2024), and Mistral-7B-Instruct-v0.3 (Jiang et al., 2023); for Socrates, we restrict evaluation to Gemma-2-27B given its increased reasoning difficulty. We use Qwen3-32B to extract bridge objects from self-NLEs and Patchscopes (Ghandeharioun et al., 2024) as interpreter. We report Latent Hop 1 Correctness, Self-NLE Faithfulness, and Self-NLE Correctness, stratified by prediction accuracy. Details on circuit $\Gamma$ and prompts to extract the bridge object and run Patchscopes are provided in Appendices D.1 and D.2.
Since NeuroFaith relies on Qwen3-32B to extract bridge objects, we evaluate its reliability for this task. When prompted to retrieve bridge objects from complete 2-hop reasoning chains ($o_{1}$, $r_{1}$, $o_{2}$) and ($o_{2}$, $r_{2}$, $o_{3}$), the latter achieves 99.5% accuracy. We also report results using Phi-4 (Abdin et al., 2024) as an alternative extractor in AppendixD.2, Table 7, showing approximately 90% correlation with Qwen3-32B. These findings confirm that NeuroFaith is robust to the choice of auxiliary model and concepts are reliably extracted.
@@
Table 1 shows the aggregated experimental results obtained by applying NeuroFaith on Wikidata-2-hop. All the analyzed metrics are higher on average for accurate predictions and improve with model size. These results support that correct reasoning tend to produce more faithful and correct self-NLE. Key failure patterns emerge with Gemma: approximately 45% of inaccurate predictions correctly resolve the first hop operation, indicating that failure often occurs in the second reasoning step ($o_{2},r_{2},o_{3}$). This suggests that models can identify correct bridge objects even when failing to complete the full reasoning chain. The difference between latent hop 1 correctness and self-NLE correctness shows that models sometimes identify correct bridge objects internally but fail to express them in their self-NLE. This gap decreases with model size, suggesting better alignment between internal and explicit reasoning in larger models. Moreover, self-NLE correctness and faithfulness are comparable between Mistral and Gemma for accurate predictions but substantially lower for Mistral on inaccurate ones.
Remarkably, even when models produce correct final answers, they identify the correct bridge object in their internal representations only 48-65% of the time. This suggests that models can arrive at correct answers through alternative reasoning pathways that bypass explicit bridge object computation. The models may be leveraging (1) direct associations between source and target entities (shortcut learning (Geirhos et al., 2020)), or (2) alternative pathways making the model accurately answer for a reason different than the ground truth bridge object (see category 9 in Figure 28). Figure 3 shows the results obtained by applying NeuroFaith on Socrates, focusing on predictions with accurate latent first hop reasoning. Like the Wikidata-2-hop dataset, accurate predictions lead to more faithful and accurate self-NLE. Table 8 in Appendix D.2 shows the complete results obtained from the Socrates dataset.
We additionally compare NeuroFaith to two commonly used faithfulness measures: CI (Atanasova et al., 2023) and AA (Wiegreffe et al., 2021). For this comparison, we use an In-Context Editing evaluation protocol, extending the approach proposed by Zaman and Srivastava (2025). Results in Appendix F (Table 16) show that NeuroFaith outperforms these baselines in 6 out of 9 settings. Finally, Appendix D.2 presents a sensitivity analysis varying circuit size. Figures 13-15 show that faithfulness scores remain consistent across circuit sizes, with only a 10% gap between single-node and full circuits, demonstrating robustness to circuit specification.
Figure 3: NeuroFaith self-NLE correctness and faithfulness on Socrates on instances with accurate first hop resolution.
@@ 5 The Case of Classification
This section instantiates NeuroFaith for classification by employing concept importance methods as mechanistic interpreter I. Given an input text $x\in\textbf{X}$, we denote $f(x)=\hat{y}\in\mathcal{Y}$ with probability score $p_{\hat{y}}(x)$ where $\mathcal{Y}$ is the label space. We assume access to a predefined set of task-relevant concepts $\mathcal{C}$ with concept labels for input texts.
@@ 5.1 NeuroFaith Instantiation for Classification
To enable mechanistic interpretation of each concept $c\in\mathcal{C}$, we compute Concept Activation Vectors (CAVs) that represent concepts as directions in $f$ representation space. Following established practices in concept-based interpretability, we employ the mean difference (diff-mean(Rimsky et al., 2024) approach for CAV computation due to its optimal balance between concept detection accuracy and computational efficiency (Wu et al., 2025). For a concept $c\in\mathcal{C}$, a token index $k$ and a layer $\ell$, the layer-wise CAV is defined as: $\overrightarrow{c_{\ell}}=\frac{1}{\left|\textbf{X}^{+}_{c}\right|}\sum_{x\in\textbf{X}^{+}_{c}}{h^{\ell}_{k}}-\frac{1}{\left|\textbf{X}^{-}_{c}\right|}\sum_{x\in\textbf{X}^{-}_{c}}{h^{\ell}_{k}}$, where $\textbf{X}^{+}_{c}$ and $\textbf{X}^{-}_{c}$ respectively represent the sets of texts from X where the concept $c$ is present or absent and $h^{\ell}_{k}$ denotes the hidden state at layer $\ell$ and token position $k$ at the specified granularity $\mathcal{G}$. We set the token index to the final position of $x$, as this location represents $f$’s complete computational state prior to next-token generation. These CAVs allow for computing the importance of each concept using Representation Engineering techniques. Following Section 3, we use an auxiliary LLM to extract a set of concepts from $e(x)$, making sure that $\left\{c_{i}\right\}_{i=1}^{p}\subset\mathcal{C}$.
@@
For a concept $c\in\left\{c_{i}\right\}_{i=1}^{p}$, we approximate its causal influence on $f(x)$ through representation engineering and concept erasure (Belrose et al., 2023). We perform controlled interventions by erasing $c$ from hidden states during forward propagation: $h^{\ell}_{k}\leftarrow h^{\ell}_{k}-\lambda\times\overrightarrow{c_{\ell}}$, where $\lambda\in[0,1]$ represents intervention intensity and with $\lambda=1$ represents maximum intervention intensity to avoid $f$ collapse (Rimsky et al., 2024). This approach provides strong causal evidence for concept importance without requiring computationally costly gradient computations such as TCAV. We define $\Gamma$ with granularity $\mathcal{G}$ set to residual stream (RS), token index as the final position (likewise CAVs), and layers are selected based on concept detectability (F1 score >> 60% using layer-wise linear probes with diff-mean). Applying this intervention across circuit $\Gamma$, we measure the resulting probability score: $p_{\hat{y}}(x,\left\{\text{do}(H^{\ell}_{k}=h^{\ell}_{k}-\lambda\times\overrightarrow{c_{\ell}})\right\}_{(k,\ell)\in\Gamma})$, where the $\text{do}(X=x)$ operator (Pearl, 2009) represents an intervention that sets variable $X$ to value $x$. We model this relationship as linear: $p_{\hat{y}}(\cdot)=\beta_{0}+\beta_{1}\times\lambda$, where the concept importance score is $\texttt{I}_{\Gamma}(c)=\beta_{1}$ (i.e. the marginal effect of intervention on probability score). $\texttt{I}_{\Gamma}(c)$ is set to 0 when its significance $t$-test shows $p>0.01$. This process is repeated for each extracted concept from $e(x)$ to derive faithfulness scores as described in Section 3.3, with $F(x,e)\in\left[0,1\right]$ for classification tasks. Details and examples are provided in Appendices D.2 and G.
Table 2: Classification accuracy and self-NLE faithfulness from NeuroFaith. ”(In)accurate represents (in)correct predictions.
Model Task Accuracy Self-NLE Faithfulness
@@
mistral-3-7B 80.6% 82.9% 86.6% 84.1% 94.8% 84.4%
@@ 5.2 Classification Experimental Analysis
We evaluate NeuroFaith’s classification instantiation on two datasets with varying complexity: AGNews (Gulli, 2005), a newspaper article classification and Ledgar (Tuggener et al., 2020), a more challenging critical domain legal document classification dataset. We apply NeuroFaith to Gemma-2 (2B, 9B, 27B) (Gemma Team, 2024) and Mistral-7B-Instruct-v0.3 (Jiang et al., 2023). We use the concept set $\mathcal{C}$ and labels of AGNews and Ledgar from Bhan et al. (2025) to compute the CAVs. We use Qwen3-32B to extract concepts from the self-NLE, ensuring extracted concepts belong to $\mathcal{C}$ for each dataset. We focus on instances solely containing concepts that are linearly detectable in at least one layer. We evaluate Self-NLE Faithfulness according to the status of the prediction preceding the self-NLE (accurate vs. inaccurate).
@@
Table 2 shows experimental results for AGNews and Ledgar datasets. Higher average self-NLE faithfulness is observed with Gemma on AGNews and Mistral on Ledgar, supporting that explanation faithfulness might correlate with task accuracy. Accurate predictions generally yield more faithful explanations, suggesting that correct predictions might rely on relevant concepts. Notably, gemma-2-9B and Mistral-3-7B respectively achieve the highest faithfulness scores on AGNews and Ledgar, going against the idea of a monotonic relationship between model scale and explanation faithfulness. Figures 8 and 10 in Appendix E.1 present examples of concepts frequently associated with unfaithful self-NLEs, revealing confusion where the model invokes concepts related to ”business” when predicting the ”world” class, and vice versa.
Comparing findings across classification and 2-hop reasoning reveals that accurate predictions tend to produce more faithful explanations, whereas scaling effects vary between tasks and models, as shown in prior work (Atanasova et al., 2023; Parcalabescu and Frank, 2024; Matton et al., 2025).
@@ 6 Linearly Detecting and Improving Self-NLE Faithfulness
Recent work has demonstrated that various safety behaviors naturally associated with faithfulness are encoded as linear directions in LLM representation spaces (Bereska and Gavves, 2024), such as hallucination (Rimsky et al., 2024) or deceptiveness (Goldowsky-Dill et al., 2025). We demonstrate that NeuroFaithfaithfulness also exhibits linear structure for (1) accurate detection in representation space and (2) manipulation to improve faithfulness through activation steering.
@@ 6.1 Linear Detection of Faithfulness
Figure 4: Faithfulness linear probe performance per model/ dataset.
@@
accurate inaccurate overall accurate inaccurate overall accurate inaccurate overall
@@
Faithfulness Linear Representation Computation.
We construct datasets of faithful and unfaithful self-NLE pairs using NeuroFaith scores. Unlike 2-hop reasoning tasks, classification self-NLE faithfulness yields continuous scores between 0 and 1. To ensure clear class separation, we focus on polarized cases where $F(x,e)\in\left\{0,1\right\}$ for clear separation. We construct input sequences as $x_{nle}=[x,f(x),e(x)]$ and extract hidden states from final token positions across all layers. We train diff-mean linear probes on these states, yielding layer-wise faithfulness representations $\overrightarrow{F_{\ell}}$. For classification tasks, we use class-averaged faithfulness vectors to isolate faithfulness-specific representations independent of class confounders. Based on these faithfulness vectors, we compute a majority vote across layer-wise linear probes beyond layer 5. Figure 4 shows faithfulness can be reliably detected across all evaluated tasks and models, with F1 scores ranging from 61.1% to 74.5%. This consistency suggests that NeuroFaith captures a systematic aspect of model behavior that manifests similarly across different tasks. Details about class-averaged faithfulness vectors and layer-wise classification score are in Appendix D.3.
Table 4: Linear vectors max. cosine similarity on gemma-2-27b with related layer. * indicates statistical significance, with $p<1\%$
AGNews Ledgar 2 hop Halluc.
[…]
Linear Representation Similarity Analysis.
To investigate the hypothesized relationships between faithfulness and related safety behaviors, we analyze the similarity between NeuroFaith faithfulness vectors and established linear representations of hallucination and deceptiveness from  Rimsky et al. (2024) and Goldowsky-Dill et al. (2025). Given that faithful self-NLE should reflect model reasoning while hallucination and deceptiveness involve misrepresentation of information, we expect their neural representations to exhibit inverse correlations. We focus on layers with F1 >> 60% where faithfulness is properly linearly represented. Table 4 shows maximum cosine similarity and corresponding layers for gemma-2-27B (additional results in Appendix D.3). Faithfulness vectors are in general negatively correlated with hallucination and deceptiveness ones (up to -0.38), aligning with expectations about the opposition between faithfulness and both deceptiveness and hallucination. Cross-task faithfulness vectors are overall positively correlated (up to 0.49), suggesting shared underlying representations of faithful across tasks. These correlations validate that NeuroFaith captures LLMs behavioral aspects aligned with established safety behaviors. These results are corroborated by Table 9 in Appendix D.3, showing cross-task transfer with gemma-2-27b: faithfulness vectors from 2-hop reasoning can detect unfaithful self-NLE in AGNews, and AGNews faithfulness vectors can detect unfaithful self-NLE in Ledgar.
6.2 Faithfulness Enhancement Through Steering on 2-hop reasoning
Having established the linear structure of NeuroFaith faithfulness representations and their negative correlation with hallucination and deceptiveness vectors, we investigate whether linear steering can improve self-NLE faithfulness during inference. We implement activation steering by modifying hidden states during inference: $h^{\ell}_{k}\leftarrow h^{\ell}_{k}+\lambda\times\overrightarrow{SV_{\ell}}$, where $\overrightarrow{SV_{\ell}}$ is the steering vector and $\lambda$ controls intervention intensity. Our objective is here to demonstrate immediate practical value for improving self-NLE faithfulness during inference without model modification. More sophisticated steering methods (Hedström et al., 2025; Vogels et al., 2025) would likely yield superior results. We evaluate three approaches: faithfulness amplification ($\overrightarrow{SV_{\ell}}=\overrightarrow{F_{\ell}}$, $\lambda=1$) and hallucination/deceptiveness inhibition ($\overrightarrow{SV_{\ell}}=\overrightarrow{H_{\ell}}$ or $\overrightarrow{D_{\ell}}$, $\lambda=-1$), where $\overrightarrow{H_{\ell}}$ and $\overrightarrow{D_{\ell}}$ are hallucination and deceptiveness linear vectors obtained from (Rimsky et al., 2024) and (Goldowsky-Dill et al., 2025). Faithfulness intervention targets only layers with F1 >> 60%, ensuring modifications occur where faithfulness is reliably encoded.
Table 3 shows steering interventions successfully convert 8-11% of unfaithful self-NLE into faithful explanations across all models. Direct faithfulness amplification slightly outperforms hallucination inhibition and substantially outperforms deceptiveness inhibition. Steering impact on faithfulness remains consistent regardless of prediction accuracy, demonstrating practical utility of NeuroFaith for real-time faithfulness enhancement without model modification. Figure 1 shows how a faithfulness linear probe identifies bridge objects as key tokens determining self-NLE faithfulness, comparing an unfaithful example with its steered faithful version. Additional analyses include comprehensive faithfulness status changes and other examples of converted self-NLE (see Appendix E.1 and H).